# Small Open Models Achieve Near-Parity with Large Models in Low-Resource Literary Translation at a Fraction of the Cost

### v1

@@ 3.1 Stage 1: Evaluating Models & Selecting Translators
The first stage identifies (i) a high-quality translation system to serve as a reference translator and (ii) a competitive open-source backbone suitable for fine-tuning. To this end, we evaluated a diverse pool of candidate systems, including commercial MT APIs, proprietary LLMs, and open-source instruction-tuned models, to gauge their suitability for the literary translation domain.
Evaluation Methodology. Because standard automatic metrics such as BLEU are poorly correlated with literary adequacy and stylistic fidelity, we relied exclusively on LLM-based human-like evaluation. Each candidate was scored on five dimensions derived from professional translation guidelines:
@@
    Accuracy: Fidelity to the full semantic content of the source.
@@
    Cultural/Pragmatic Adaptation: Adequacy of cultural references and moral framing for the target audience.
Each dimension is scored on a 1 to 5 scale (1 = poor, 5 = excellent). To conduct this evaluation at scale, we utilize an LLM as an automated judge. This model was prompted with detailed instructions to emulate a professional translation reviewer. The prompt (in natural language) set the context: “You are a professional translation evaluator. You will be given an English fable and a Romanian translation. Evaluate the translation for accuracy, fluency, coherence, style, and cultural/pragmatic fidelity. Provide a score from 1 to 5 for each category along with a brief justification. Output your evaluation in valid JSON format with fields for each score and justification.”
By constraining the output to JSON, we ensured the evaluator’s responses could be easily parsed and aggregated. We evaluated a randomly selected subset of 100 test instances with this LLM-based evaluator for each model’s translations (100 unique fables per model). The result is a set of scores across the five dimensions, as well as qualitative explanations (which we used to sanity-check the scores). Using an LLM as a judge in this manner draws on recent work such as G-Eval and GPTScore, which have shown that LLMs can approximate human evaluation of text generation to a useful degree [12]. The average scores from these evaluations are reported in Table 1, providing a multi-dimensional comparison of translation quality.
All evaluation prompts, along with example JSON outputs, are included in our repository for transparency. The JSON output for each evaluated sample contains keys like "accuracy": 5, "fluency": 4, and so forth. We emphasize that these LLM-generated scores are used as an auxiliary signal, not as an absolute ground truth; nonetheless, they offer valuable insight into where certain models might be failing (e.g., lower style scores despite strong accuracy).
Selecting Translators and Backbones. From these evaluations, the system with the highest rubric scores was designated as the reference translator for corpus creation and benchmarking. Separately, the strongest-performing open-source model was selected as the backbone for fine-tuning in Stage 3. This dual selection ensures both high-quality references for evaluation and open, reproducible foundations for model adaptation.
@@ 3.2 Stage 2: Parallel Corpus Generation (15k Fables)
A central step in constructing a robust benchmark for low-resource literary MT is the creation of a high-quality parallel dataset tailored to the narrative and stylistic demands of the target genre. In TF2, we address the acute shortage of English–Romanian literary data by leveraging state-of-the-art LLMs to synthesize a sizable and diverse set of story translations, adhering to the best practices in synthetic corpus construction [14, 4].
The pipeline begins by sampling 15,000 short fables from the DS-TF1-EN-3M dataset [^6] https://huggingface.co/datasets/klusai/ds-tf1-en-3m, each consisting of a concise narrative culminating in an explicit moral. Rather than relying on scarce human-translated sources, we use the top-performing LLM-based translator, selected via rigorous benchmarking, to generate Romanian counterparts for each English story. This process yields a high-coverage, machine-generated parallel corpus, designed to serve as both a training resource and a reference benchmark for literary MT.
The resulting TinyFabulist DS-TF2-EN-RO-15K dataset is released on Hugging Face [^7] https://huggingface.co/datasets/klusai/tf2-en-ro-15k under the permissive MIT License, ensuring full transparency and broad community reuse. Each record is stored in JSONL (one JSON object per line) and includes the English source, the Romanian target, and comprehensive metadata to facilitate downstream analysis and reproducibility. Key fields are:
  • fable: The original English fable text, typically 1–3 paragraphs, always ending with an explicit moral.
@@
  • pipeline_stage: Stage of the data-generation pipeline that produced the record (e.g., translation).
@@
  • prompt_hash: SHA-256 hash of the generation prompt, used for deduplication and reproducibility.
  • llm_name: Identifier of the model used for translation (path or Hugging Face snapshot).
  • translation_model: Duplicate field pointing to the model checkpoint that produced the translation, for explicit traceability.
  • generation_timestamp: Unix timestamp of when the translation was generated, enabling chronological auditing.
The dataset is split into 12,000 training pairs, 1,500 validation pairs, and 1,500 test pairs, supporting robust supervised learning and benchmarking. Although the Romanian references are model-generated, their quality is ensured by selecting the strongest available translator through the quality-aware benchmarking pipeline outlined in Section 3.1. This pragmatic approach is aligned with recent best practices for corpus creation in low-resource settings, and each entry’s rich metadata supports reproducibility, cross-evaluation, and longitudinal studies.
Overall, the TinyFabulist DS-TF2-EN-RO-15K corpus provides a transparent, extensible foundation for training, evaluation, and methodological innovation in literary machine translation, directly addressing the lack of large-scale, genre-specific resources for English–Romanian and similar language pairs.
@@ 3.3 Stage 3: Fine-Tuning—Parameter-Efficient Domain Adaptation
The third stage adapts open instruction-tuned LLMs to the literary translation domain using parameter-efficient fine-tuning. Instead of updating all model parameters, which is costly and less reproducible, we employ Low-Rank Adaptation (LoRA; 7), a technique that injects lightweight, trainable adapters into attention and feed-forward layers while keeping most parameters frozen. This approach yields strong domain adaptation with minimal computational overhead.
@@
    Tokenization and masking: Employ the native SentencePiece or BPE vocabulary associated with each model family. In the labels tensor, mask all tokens except the Romanian output, ensuring that only target-side predictions contribute to the training loss. English inputs are tokenized but ignored in the loss function; only target tokens contribute to gradients.
@@
    Adapter injection: We attach LoRA adapters to all major projection matrices in both self-attention and feed-forward blocks—q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj. For each weight $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ we learn a low-rank update $\Delta W=\tfrac{\alpha}{r}BA$ with r=32r{=}32, $\alpha{=}32$ (effective scaling =1.0=1.0), and apply dropout p=0.05p{=}0.05 on the adapter path. We train only $A,B$ while freezing the backbone (no bias adaptation), and merge $\Delta W$ into $W$ at inference. This follows best practices in recent literature [15, 3].
@@
    Training regime: Fine-tune models using AdamW optimization, gradient accumulation, cosine learning rate scheduling, and mixed-precision (FP16 or bfloat16) arithmetic. To maximize sample efficiency and prevent overfitting, we combine LoRA dropout with early stopping based on held-out validation loss, following Liu et al. [12].
@@
    Deployment: After training convergence, we merge LoRA adapters into the base weights for standalone inference and produce 8-bit quantized variants [2, 5] to further reduce memory and latency overhead. We additionally leverage llmcompressor to generate W8A8 quantized models, which integrate seamlessly with vLLM for efficient inference [10, 23]. For offline distribution we also export GGUF artifacts compatible with llama.cpp and Hugging Face GGUF endpoints888 https://github.com/ggml-org/llama.cpp.
Evaluation. Fine-tuned models are assessed using both rubric-based evaluation (as in Stage 1) and BLEU. Here, BLEU is used as a complementary metric by comparing translations against the reference corpus created in Stage 2. While rubric scores remain the primary measure of literary adequacy, BLEU provides a lightweight consistency check on lexical overlap, complementing rubric scores without serving as a stand-alone measure.
Backbone Coverage. This fine-tuning procedure is applied across multiple model scales, from compact to large, to demonstrate robustness of domain adaptation under varying computational constraints. Results are reported in Section 4, highlighting improvements in both rubric scores and BLEU relative to untuned backbones.
@@ 3.4 Stage 4: Large-Scale Fable Translation
The final stage of the framework is the large-scale translation of the entire TinyFabulist English corpus into Romanian, leveraging our fine-tuned models for efficient and high-quality generation. DS-TF1-EN-3M comprises approximately 3 million AI-generated English fables, all of which were systematically translated into Romanian using our best-performing system, as well as other fine-tuned checkpoints where appropriate. This process produced the DS-TF2-EN-RO-3M dataset: a massive, openly available English–Romanian parallel corpus specifically tailored for literary translation research.
Thanks to the efficiency of the TF2 models, the translation of millions of fables was completed on modest hardware (e.g., small GPU clusters or standard cloud infrastructure) without prohibitive computational cost. The resulting dataset, which vastly surpasses the initial 15k seed corpus in both size and diversity, is now released for the research community and accessible on the Hugging Face Hub. [^9] https://huggingface.co/datasets/klusai/tf2-en-ro-3m
A full statistical and structural overview of the DS-TF2-EN-RO-3M dataset—including metadata schema, field definitions, and example use cases—is provided in Section 5. This large-scale resource supports new directions in literary MT, cross-lingual narrative analysis, and low-resource NLP, furthering the “living corpus” philosophy of TinyFabulist.
@@ 4 Experiments and Results
This section evaluates TF2 ’s translation models against a wide array of strong baselines—both proprietary and open—using both automated and LLM-based metrics. We present a comprehensive comparison of all fine-tuned KlusAI models, their quantized variants, and the influence of decoding temperature on translation quality.
@@ 4.1 Benchmarked Systems and Baselines
Table 1 shows that GPT-o3 attains near-ceiling values across all rubric dimensions (4.86–4.97), surpassing the next-best systems (e.g., GPT-4.1, Gemini-2.5) by a consistent margin. We therefore adopt GPT-o3 outputs as our silver-standard references for automatic comparisons elsewhere in the paper. Importantly, subsequent BLEU numbers reported for other systems use these GPT-o3 translations as references, which explains why GPT-o3 itself does not appear with a comparable BLEU entry in those tables.
@@
Among open baselines, Gemma-3-12B-it is the strongest model on our rubric (Avg. 4.43), clearly outperforming the 4B (3.81) and 1B (2.02) variants. It also provides a permissive license and robust tooling, making it a practical and competitive foundation for adaptation. Consequently, we select Gemma-3-12B-it (and, for scaling experiments, its 4B and 1B counterparts) as the backbones for the TF2 fine-tuning in the next subsection.
@@ 4.2 Overall Translation Quality with TF2 Models
The TF2 series comprises three parameter-efficiently fine-tuned variants of the Gemma-3 instruction-tuned backbones: Gemma-3-1B-it, Gemma-3-4B-it, and Gemma-3-12B-it. Each model was adapted on the DS-TF2-EN-RO-15K corpus using LoRA adapters injected into all major projection matrices (q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj) with configuration r=32r{=}32, $\alpha{=}32$, and dropout p=0.05p{=}0.05. Training followed standard low-rank adaptation practice [15], with adapters merged into the base weights after convergence. For deployment, we additionally provide 8-bit quantized variants, as well as a distilled TF2-1B checkpoint derived from teacher–student compression. This setup yields three scalable, self-hostable translation models—TF2-1B, TF2-4B, and TF2-12B—covering the small, medium, and large open-model parameter ranges. Building upon these strong, diverse baselines, we next evaluate all three fine-tuned TF2 models—TF2-1B, TF2-4B, TF2-12B—each directly compared to its original backbone: Gemma-3-1B-it, Gemma-3-4B-it, and Gemma-3-12B-it, respectively. This strategy enables a direct and transparent assessment of the gains achieved through parameter-efficient adaptation on domain-specific, synthetic literary data.
All three TF2 models show dramatic improvements over their untuned checkpoints, closing much of the gap to the top-tier proprietary LLMs. To probe decoding robustness, we evaluate each model under different temperature ($T$) settings: $T=0.0$ (deterministic greedy decoding), $T=0.2$ (low-variance sampling), and $T=1.0$ (higher-variance sampling). Below, we summarize the key results for each parameter scale.
@@
  • TF2-1B (Gemma-3-1B-it backbone): The smallest model, after fine-tuning, attains an average rubric score of 3.75, representing a significant jump from its untuned base (2.02). BLEU, however, remains modest (0.2180 for the TF2-1B-distilled variant; 0.0543 for the standard TF2-1B checkpoint), reflecting the inherent limitations of compact architectures for long-form narrative translation.
  • TF2-4B (Gemma-3-4B-it backbone): The 4B parameter model (both FP16 and 8-bit quantized) consistently achieves averages above 4.65 across all five evaluation dimensions, peaking at 4.74 (T=0.0). BLEU ranges from 0.0496 (quantized, T=0.2) to 0.1290 (quantized, T=0.0) and 0.1278 (FP16, T=1.0), highlighting strong quantization and decoding robustness. This represents a substantial qualitative improvement over the untuned Gemma-3-4B-it base (avg. rubric 3.81, BLEU 0.1005).
  • TF2-12B (Gemma-3-12B-it backbone): The largest model in our suite delivers the strongest results among all open, self-hostable systems. At optimal decoding ($T=0.0$), TF2-12B attains an average rubric score of 4.83 and BLEU 0.0926; at $T=0.2$, 4.82 and BLEU 0.0647; even at $T=1.0$, it maintains a robust 4.66 (BLEU 0.0784). The 8-bit quantized version is nearly indistinguishable, , in line with prior findings on activation-aware quantization [11], showing that aggressive quantization incurs negligible quality loss. All these numbers represent clear, multi-dimensional gains over the untuned Gemma-3-12B-it base (avg. rubric 4.43, BLEU 0.0214).
@@
A summary of main results—with both BLEU and LLM-based rubric scores—is provided in Tables 2 and 3. For each model scale, we report the fine-tuned TF2 model, its quantized variant (where applicable), and the corresponding Gemma backbone for direct comparison.
@@
Table 2: BLEU scores for each fine-tuned TF2 model and its corresponding Gemma backbone. Quantized and distilled variants are included. BLEU reflects test set n-gram overlap with the GPT-o3 reference and is reported on a normalized 011 scale ($0.0926\equiv 9.26$ BLEU points on the conventional scale). Lower BLEU for Gemma-3-12B-it vs Gemma-3-4B-it arises from greater paraphrasing and structural divergence by the larger model, which BLEU penalizes despite higher adequacy and fluency confirmed by rubric scores.

### v2

@@ 3.1 Stage 1: Evaluating Models & Selecting Translators
The first stage serves two objectives: (i) selecting a translation system to act as a reference translator for parallel corpus construction, and (ii) identifying a competitive open-source backbone suitable for downstream parameter-efficient fine-tuning. We evaluate a diverse pool of candidate systems, including commercial MT APIs, proprietary LLMs, and open instruction-tuned models, with a focus on English–Romanian literary translation.
Evaluation Methodology. Translation quality in literary narratives involves multiple complementary aspects, including semantic fidelity, linguistic well-formedness, discourse coherence, stylistic appropriateness, and cultural adaptation. To capture these facets in a scalable and comparable manner, we employ rubric-based LLM evaluation at Stage 1 and introduce reference-based metrics (e.g., BLEU) once a consistent reference corpus becomes available in later stages. Targeted human assessment is used to anchor the interpretation of automated scores (Section 4.5).
At Stage 1, where no human-authored Romanian references are available, candidate systems are compared primarily using a multi-dimensional rubric implemented via an LLM acting as an automated judge. This rubric evaluates five dimensions derived from professional translation guidelines:
@@
    Accuracy: Fidelity to the semantic content of the source text.
@@
    Cultural/Pragmatic Adaptation: Appropriateness of cultural references and moral framing for the target audience.
Each dimension is scored on a 1–5 Likert scale (1 = poor, 5 = excellent). To operationalize this rubric at scale, we prompt an LLM to emulate a professional literary translation reviewer:
“You are a professional translation evaluator. You will be given an English fable and a Romanian translation. Evaluate the translation for accuracy, fluency, coherence, style, and cultural/pragmatic fidelity. Provide a score from 1 to 5 for each category, along with a brief justification. Output your evaluation in valid JSON format with fields for each score and justification.”
The structured JSON output enables reliable aggregation of scores across systems. For each candidate model, we evaluate a randomly sampled subset of 100 distinct fables, yielding per-dimension scores and short qualitative rationales. The resulting averages provide a comparative view of translation behavior across systems. This approach follows prior work showing that LLM-based rubrics can provide useful comparative signals in text generation evaluation [12]. In our pipeline, these scores are used as an auxiliary decision signal for model selection rather than as an absolute measure of translation quality.
Selecting Translators and Backbones. Based on aggregated rubric scores, the highest-ranked system is selected as the reference translator for parallel corpus construction and downstream benchmarking. Independently, the strongest-performing open-source model is selected as the backbone for parameter-efficient fine-tuning in Stage 3.
@@ 3.2 Stage 2: Parallel Corpus Generation (15k Fables)
A central step in constructing a benchmark for low-resource literary MT is the creation of a parallel dataset tailored to the narrative and stylistic demands of the target genre. In TF2, we address the shortage of English–Romanian literary data by leveraging LLMs to generate a diverse set of story translations, following established practices in synthetic corpus construction [14, 4].
We sample 15,000 short fables from DS-TF1-EN-3M  [^6] https://huggingface.co/datasets/klusai/ds-tf1-en-3m, each consisting of a concise narrative ending with an explicit moral. We then translate them into Romanian using the top-ranked translator selected in Stage 1. This yields a machine-generated parallel corpus intended to serve as both a training resource and a consistent reference set for subsequent benchmarking.
The resulting TinyFabulist DS-TF2-EN-RO-15K dataset is released on Hugging Face [^7] https://huggingface.co/datasets/klusai/tf2-en-ro-15k under the permissive MIT License. Each record is stored in JSONL format and includes the English source, the Romanian translation, and metadata to support downstream analysis and reproducibility. Key fields include:
  • fable: The original English fable text (typically 1–3 paragraphs), ending with an explicit moral.
@@
  • pipeline_stage: Stage of the data-generation pipeline that produced the record (here translation).
@@
  • prompt_hash: SHA-256 hash of the translation prompt, used for reproducibility and traceability.
  • llm_name: Identifier of the model used for translation (e.g., model name or Hugging Face snapshot).
  • translation_model: Identifier of the specific checkpoint used to produce the translation (when distinct from llm_name).
  • generation_timestamp: Unix timestamp indicating when the translation was generated.
The dataset is split into 12,000 training pairs, 1,500 validation pairs, and 1,500 test pairs. Although the Romanian references are model-generated, their selection follows the benchmarking procedure described in Section 3.1. The accompanying metadata supports reproducibility, comparative evaluation, and auditing of translation provenance.
Overall, TinyFabulist DS-TF2-EN-RO-15K provides a transparent foundation for training and evaluating English–Romanian literary MT models in the absence of large human-translated corpora.
@@ 3.3 Stage 3: Fine-Tuning—Parameter-Efficient Domain Adaptation
The third stage adapts open instruction-tuned LLMs to the literary translation domain using parameter-efficient fine-tuning. Instead of updating all model parameters, we employ Low-Rank Adaptation (LoRA; 7), which injects lightweight, trainable adapters into attention and feed-forward layers while keeping most parameters frozen. This approach enables domain adaptation with relatively modest computational overhead.
@@
    Tokenization and masking: Employ the native SentencePiece or BPE vocabulary associated with each model family. In the labels tensor, mask all tokens except the Romanian output, ensuring that only target-side predictions contribute to the training loss.
@@
    Adapter injection: Attach LoRA adapters to major projection matrices in both self-attention and feed-forward blocks—q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj. For each weight $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ we learn a low-rank update $\Delta W=\tfrac{\alpha}{r}BA$ with r=32r{=}32, $\alpha{=}32$ (effective scaling =1.0=1.0), and apply dropout p=0.05p{=}0.05 on the adapter path. We train only $A,B$ while freezing the backbone (no bias adaptation), and merge $\Delta W$ into $W$ at inference. This follows best practices in recent literature [15, 3].
@@
    Training regime: Fine-tune models using AdamW optimization, gradient accumulation, cosine learning rate scheduling, and mixed-precision (FP16 or bfloat16) arithmetic. To reduce overfitting, we apply early stopping based on held-out validation loss.
@@
    Deployment: After training convergence, merge LoRA adapters into the base weights for standalone inference and produce 8-bit quantized variants [2, 5] to reduce memory and latency. We additionally leverage llmcompressor to generate W8A8 quantized models, compatible with vLLM [10, 23]. For offline distribution we also export GGUF artifacts compatible with llama.cpp and Hugging Face GGUF endpoints [^8] https://github.com/ggml-org/llama.cpp.
Evaluation. Fine-tuned models are assessed using both rubric-based evaluation (as in Stage 1) and BLEU. Here, BLEU is used as a complementary metric by comparing translations against the reference corpus created in Stage 2. Rubric scores remain the primary measure for multi-dimensional translation quality, while BLEU provides a lightweight consistency check on lexical overlap.
Backbone Coverage. This fine-tuning procedure is applied across multiple model scales to assess robustness under varying computational constraints. Results are reported in Section 4, including changes in both rubric scores and BLEU relative to untuned backbones.
@@ 3.4 Stage 4: Large-Scale Fable Translation
The final stage translates the full TinyFabulist English corpus into Romanian using the fine-tuned models. DS-TF1-EN-3M comprises approximately 3 million AI-generated English fables, which we translate into Romanian primarily using the best-performing fine-tuned checkpoint. In addition, for a small number of runs we also generate translations with alternative fine-tuned checkpoints to support ablations and comparative analyses reported in Section 4. This process yields the DS-TF2-EN-RO-3M dataset, an openly released English–Romanian parallel corpus for literary translation research.
The translation process is designed to be feasible on modest hardware (e.g., small GPU clusters or standard cloud infrastructure) without prohibitive computational cost. The resulting dataset is released on the Hugging Face Hub. [^9] https://huggingface.co/datasets/klusai/tf2-en-ro-3m
A full statistical and structural overview of the DS-TF2-EN-RO-3M dataset—including metadata schema, field definitions, and example use cases—is provided in Section 5.
@@ 4 Experiments and Results
This section evaluates TF2 ’s translation models against a range of strong baselines—both proprietary and open—using reference-based metrics (BLEU), an LLM-based rubric, and a targeted human evaluation. We compare all fine-tuned TF2 models, their quantized variants, and the influence of decoding temperature on translation quality.
@@ 4.1 Benchmarked Systems and Baselines
Table 1 indicates that GPT-o3 attains the highest average rubric score across the evaluated systems. We therefore adopt GPT-o3 outputs as silver-standard references for reference-based comparisons elsewhere in the paper. Subsequent BLEU results use these GPT-o3 translations as references, which is why GPT-o3 itself does not appear with a comparable BLEU entry in those tables.
@@
Among the open baselines, Gemma-3-12B-it achieves the highest rubric average (4.43), outperforming the 4B (3.81) and 1B (2.02) variants. It also offers a permissive license and mature tooling, making it a practical foundation for adaptation. We therefore select Gemma-3-12B-it (and, for scaling experiments, its 4B and 1B counterparts) as backbones for TF2 fine-tuning.
@@ 4.2 Overall Translation Quality with TF2 Models
The TF2 series comprises three parameter-efficient fine-tuned variants of the Gemma-3 instruction-tuned backbones: Gemma-3-1B-it, Gemma-3-4B-it, and Gemma-3-12B-it. Each model is adapted on the DS-TF2-EN-RO-15K corpus using LoRA adapters injected into all major projection matrices (q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj) with configuration r=32r{=}32, $\alpha{=}32$, and dropout p=0.05p{=}0.05. Training follows standard low-rank adaptation practice [15], with adapters merged into the base weights after convergence. For deployment, we provide 8-bit quantized variants, as well as a distilled TF2-1B checkpoint derived from teacher–student compression. This yields three self-hostable translation models—TF2-1B, TF2-4B, and TF2-12B—covering small, medium, and large parameter ranges.
We compare each fine-tuned TF2 model directly to its corresponding untuned backbone: Gemma-3-1B-it, Gemma-3-4B-it, and Gemma-3-12B-it. We also evaluate decoding robustness across temperatures $T\in\{0.0,0.2,1.0\}$.
We observe that BLEU scores and rubric-based evaluations do not always correlate monotonically across models and decoding settings. This divergence is expected in literary translation, where higher-quality outputs may involve paraphrasing, restructuring, and stylistic variation that reduce n-gram overlap with a single reference. Accordingly, we interpret BLEU primarily as a consistency signal relative to the chosen reference translation, and not as a proxy for overall literary quality.
Below, we summarize the key results for each parameter scale.
@@
  • TF2-1B (Gemma-3-1B-it backbone): After fine-tuning, the smallest model attains an average rubric score of 3.75 (up from 2.02). BLEU remains comparatively low (0.2180 for TF2-1B-distilled; 0.0543 for TF2-1B), consistent with the difficulty of long-form narrative translation for compact architectures.
  • TF2-4B (Gemma-3-4B-it backbone): The 4B model achieves rubric averages above 4.5 across all temperatures, peaking at 4.74 ($T=0.0$). BLEU ranges from 0.0496 (quantized, $T=0.2$) to 0.1290 (quantized, $T=0.0$) and 0.1278 (FP16, $T=1.0$). Relative to the untuned backbone (avg. rubric 3.81, BLEU 0.1005), the fine-tuned variants show consistent gains under rubric-based evaluation.
  • TF2-12B (Gemma-3-12B-it backbone): The largest model provides the strongest results among the open TF2 models. At $T=0.0$, TF2-12B attains an average rubric score of 4.83 with BLEU 0.0926; at $T=0.2$, 4.82 with BLEU 0.0647; and at $T=1.0$, 4.66 with BLEU 0.0784. The 8-bit quantized version remains close in rubric scores across temperatures, consistent with prior findings on activation-aware quantization [11]. These results improve substantially over the untuned Gemma-3-12B-it baseline (avg. rubric 4.43, BLEU 0.0214).
@@
A summary of the main results—with both BLEU and LLM-based rubric scores—is provided in Tables 2 and 3. For each model scale, we report the fine-tuned TF2 model, its quantized variant (where applicable), and the corresponding Gemma backbone for direct comparison.
@@
Table 2: BLEU scores for each fine-tuned TF2 model and its corresponding Gemma backbone. Quantized and distilled variants are included. BLEU reflects test set n-gram overlap with the GPT-o3 reference and is reported on a normalized 011 scale ($0.0926\equiv 9.26$ BLEU points on the conventional scale). Lower BLEU for Gemma-3-12B-it vs. Gemma-3-4B-it can arise from greater paraphrasing and structural divergence by the larger model, which BLEU penalizes despite higher rubric scores.